OpenStudy (anonymous):

What is the solution set for |–2x + 8| = 12 x = –2 and x = 10 x = –2 and x = –10 x = 2 and x = 10 x = 2 and x = –10

OpenStudy (fifciol):

-2x+8=12 or -2x+8=-12 -2x=4 or -2x =-20 x=-2 or x=10 first answer